# Co-designing an AI Impact Assessment Report Template with AI Practitioners and AI Compliance Experts

### v1

@@ Introduction
As it has become evident that AI system impacts involve not only the underlying algorithms (Watkins et al. 2021; Barnett and Diakopoulos 2022), but also their systemic impacts (Ehsan et al. 2022; Shelby et al. 2023), there has been a growing demand for comprehensive AI impact assessments (AIIAs) (Selbst 2021). In the absence of regulatory requirements and standardized frameworks, initially, AIIAs have been regarded as an extension of existing AI governance processes, such as assessments of privacy (Wright and Wadhwa 2012), data protection (Janssen 2020), and social consequences of AI systems, models and services (Selbst 2021; Raji et al. 2020). However, over time, AIIAs developed into an independent component of AI governance (Skoric 2023). By 2021, a total of 38 distinct impact assessment methods have been introduced (Stahl et al. 2023) serving various goals such as prompting companies to proactively address social consequences (Sherman and Eisenberg 2024), aligning system behavior with organization’s responsible AI principles (Raji et al. 2020; Fujitsu Research Center for AI Ethics 2022; Microsoft 2022b), increasing awareness of potential harms among development teams (Johnson and Heidari 2023), and documenting decisions to facilitate learning and future governance development (Selbst 2021).
However, current AIIAs reports are marked by a lack of comprehensiveness and insufficient grounding in established frameworks such as the EU AI Act (European Comission 2024), the NIST’s AI Risk Management Framework (AI RMF) (National Institute of Standards and Technology 2023a), and international standards such as the ISO 42001 AI Management System (ISO/IEC 2023), that provide complementary angles for impact assessment: regulatory compliance (EU AI Act), risk management (AI RMF), and organizational best practices (ISO 42001). The lack of alignment with regulations makes it harder for stakeholders to understand the broad societal, environmental, and economic impacts of AI uses, which are essential for risk assessments (European Comission 2024; Hupont et al. 2024). Instead, they must repeatedly gather information across disparate assessments of various AI components (e.g., models, systems, and algorithms) rather than focusing on specific uses. Developer teams, in particular, often encounter difficulties in initiating AI impact assessments (Buçinca et al. 2023) and require additional guidance throughout this process (Wang et al. 2023). In response to these issues, our aim is to examine recent regulatory frameworks and existing AIIAs to devise together with AI practitioners and AI compliance experts a reporting template focusing on the intended use of the system. This template aims to provide the necessary information for conducting impact assessments based on regulations across various AI system uses and to be adaptable to different roles. In so doing, we made two contributions:
@@
    We designed a comprehensive template for an impact assessment report that is grounded in regulations, and did so by conducting two studies. First, we elicited an initial set of design requirements through literature review and semi-structured interviews with 2 AI compliance experts. Second, we co-designed the impact assessment report template with 14 AI practitioners and 6 AI compliance experts, grounded it in regulatory requirements of the EU AI Act, NIST’s AI Risk Management Framework (AI RMF), and ISO 42001 (AI Management System).
@@
    We populated the template with a real-world use case of an AI system (a companion app designed to improve the meeting experience) and evaluated it with eight AI practitioners and five AI compliance experts from academia and industry. We compared the final report to a baseline report derived from the typical structure of existing impact assessment reports. Participants found that the final report provided more complete information for impact assessments and addressed all AI system components and impacts more broadly than the baseline. Both the final report and the baseline were rated as highly adaptable to different AI uses and adaptable to different roles.
@@
Figure 1: Overview of our three-step method for designing a comprehensive template for an impact assessment report grounded in the EU AI Act, NIST’s AI RMF, and ISO 42001. In the first step, we interviewed AI compliance experts to elicit two design requirements and design the first prototype of the template. In the second step, we engaged both AI practitioners and AI experts to iteratively elicit two additional design requirements and co-design five iterations of the template. In the third step, we populated the final version of the template with AI practitioners’ responses and evaluated it against our four design requirements and the baseline template identified in the literature review of AIIAs (Stahl et al. 2023).
@@
AI impact assessments are defined as structured processes for understanding the implications of proposed AI systems (Stahl et al. 2023). They have been proposed as both law-agnostic self-regulating processes (National Institute of Standards and Technology 2023a, b) and official processes in regulations (European Comission 2024; The Danish Institute for Human Rights 2023; The Government of Canada 2023) and organizational standards (ISO/IEC 2023). For example, the NIST AI RMF, a voluntary guidance framework for organizations that design, develop, deploy, or use AI systems, suggests assessing the beneficial and harmful impacts of AI on individuals, groups, organizations, and society. The EU AI Act (European Comission 2024), the first comprehensive AI regulation in the European Union, requires comprehensive fundamental rights impact assessment for high-risk AI uses deployed by bodies governed by public law, private operators providing public services, and operators assessing creditworthiness and conducting risk assessments for life and health insurance. This impact assessment, conducted before deployment and updated as needed, should cover affected groups, risks of harm, human oversight, and risk management strategies. Similarly, the world’s first ISO 42001 standard for AI management system (ISO/IEC 2023) requires organizations to assess the potential consequences of AI systems on individuals and societies, including human rights, legal positions, and life opportunities.
Despite ongoing efforts to standardize AIIAs (ISO/IEC 2025), a consensus on their content and reporting methods for stakeholders and the public is still lacking (Watkins et al. 2021; Sherman and Eisenberg 2024; European Comission 2024). Stahl et al. (2023) identified and studied 38 AIIAs, showing that they vary widely in topics, focus, and formats, primarily covering human rights (The Council of Europe 2018; Mantelero 2022; The Danish Institute for Human Rights 2023) and ethics (Gebru et al. 2021; The High-Level Expert Group on Artificial Intelligence 2020). The focus of AIIAs varies, assessing the impacts of AI models, systems, or services, typically mentioning use cases only briefly in final reports. However, there is growing agreement that impact assessments should be conducted on specific uses (European Comission 2024; Hupont et al. 2024). In the EU AI Act (European Comission 2024), the specific use determines one of four risk categories (unacceptable, high, limited, minimal), each with its own legal requirements. Similarly, in the NIST framework (National Institute of Standards and Technology 2023a), use-case profiles are required to describe the current state and the desired state of the system, allowing for risk management at different stages of the AI lifecycle.
@@
Researchers highlight the need for more inclusive, value-sensitive co-design processes (Johnson and Heidari 2023; Sadek et al. 2024) that explicitly incorporate compliance experts (Hupont et al. 2024; Constantinides et al. 2024a). For example, OpenLoop’s policy prototyping experiment involved developers, researchers, policymakers, and regulators collaborating to create a new policy and documentation of potential harms and mitigations for automated decision-making systems (Gomes de Andrade and Kontschieder 2021). Similarly, the creators of use case cards detailing AI systems’ intended uses engaged in co-design workshops with UML experts, and EU policy experts (Hupont et al. 2024).
@@ Methods
To design a comprehensive template for an impact assessment report that is grounded in regulations, we conducted a series of three studies (Figure 1). These studies included literature review and interviews with 2 AI compliance experts, an iterative co-design process with 14 AI practitioners and 6 AI compliance experts, and a user study to evaluate the template with 8 AI practitioners working in industry research and 5 AI compliance experts from industry and academia.
@@
To elicit requirements for designing a comprehensive template for an impact assessment report that is grounded in regulations, we resorted to previous literature and did so in two steps. First, we reviewed prior academic literature investigating the use of risk assessments, impact reports, and AI documentation for compliance (Skoric 2023; Selbst 2021; Gaumond and Régis 2023; National Institute of Standards and Technology 2023a). We also drew upon five existing AI impact assessment templates (Microsoft 2022b; Fujitsu Research Center for AI Ethics 2022; National Institute of Standards and Technology 2023b; Ada Lovelace Institute 2022; The Government of Canada 2023) and value-sensitive games and cards (Ballard, Chappell, and Kennedy 2019; Vakkuri et al. 2021; Sadek et al. 2024). Second, we identified the necessary minimum information to conduct impact assessment in line with three frameworks: the EU AI Act (European Comission 2024), the NIST’s AI RMF (National Institute of Standards and Technology 2023a), and the ISO 42001 (ISO/IEC 2023). We selected these frameworks because they have undergone a consensus-driven, transparent process with wide consultations involving various roles such as developers, researchers, managers, and compliance experts. Together, they provide three complementary angles for impact assessment: the EU AI Act offers regulatory compliance, setting a precedent for other AI regulations; the AI RMF provides practical insights for risk management in corporate environments; and ISO 42001 emphasizes organizational best practices. In this step, we started from reviewing the EU AI Act (European Comission 2024) and identifying relevant excerpts that pertain to risk management (Article 9), data governance (Article 10), system monitoring (Articles 14, 15, 72), technical documentation (Articles 11, 12, 18, Annex IV), system transparency (Articles 13, 50), impact assessment (Article 27), and provider obligations (Articles 53, 55). We then reviewed the NIST’s AI RMF (National Institute of Standards and Technology 2023a) and its Algorithmic Impact Assessment template (National Institute of Standards and Technology 2023b). From these two sources, we identified relevant excerpts related to mapping the risks of a system’s use, minimizing its negative impacts while maximizing its benefits, and communicating these aspects to various stakeholders. Finally, from the ISO 42001, we identified excerpts on documenting resources for AI systems (data, tools, systems, computing, and human resources), assessing impacts on individuals, groups of individuals, and society, and reporting AI system details to stakeholders.
@@
Having these excerpts from previous literature at hand, we conducted 30-minute semi-structured interviews with 2 AI compliance experts to jointly annotate the collected excerpts (i.e., the EU AI Act, the NIST’s AI RMF, and the ISO 42001) with high-level topics. We then jointly organized the topics in a thematic and hierarchical manner, resulting in the creation of the first version of the template (Figure 2, V1). Through this joint exercise, we surfaced two main design requirements for an impact assessment report template:
@@
Figure 2: Overview of the iterative co-design process involving 14 AI practitioners and 6 AI compliance experts. Initially, we engaged only AI practitioners (P1-P4), but included other practitioners (P5-P14) and compliance experts (E3-E8) in later iterations. Individual sessions with 3 to 6 participants revealed key usability issues, prompting template iterations tested with new participants. By the fifth iteration, we had gathered sufficient information from the co-design activities, with no new significant insights or usability issues emerging. Therefore, we stopped the activities and transitioned to the template evaluation user study with AI practitioners (P15-P22) and compliance experts (E9-E13).
@@
After designing the first version of the template, we conducted a series of individual, 30-minute co-design sessions with 14 AI practitioners and 6 AI compliance experts (Figure 2). These sessions aimed to ensure that the template provided sufficient information for impact assessment according to the EU AI Act, NIST’s AI RMF, and ISO 42001, and was usable by various roles in the assessment process.
Participants. We aimed to achieve a diverse participant sample using snowball sampling, a method where existing study participants recruit future participants from among their acquaintances. We began by identifying 6 initial participants through an internal mailing list at a large tech company, asking for individuals who were familiar with the EU AI Act and had at least one ongoing AI project. These participants were then asked to refer additional participants from their networks, thus expanding the sample size through successive referrals. We recruited a total of 20 participants (13 male, 7 female, with a median age of 34 years old) representing a variety of AI practitioners such as researchers (9), engineers (2), managers (3), designers (1), and AI compliance experts (6), and potential end users of the template. Their expertise span across various areas including generative AI, machine learning, deep learning, computer vision, AI standardization, and human rights. Additionally, each participant was actively involved in developing or evaluating at least one ongoing AI project during the time of the interviews. AI practitioners offered practical examples on navigating regulatory challenges in their projects while maintaining innovation. Compliance experts contributed their knowledge of existing regulations, insights from ongoing regulatory works, and lessons learned from reviewing AI system uses. Table 1 provides an overview of the participants’ demographics and relevant experience.
Procedure. Before the session, we emailed participants with a demographic survey, a brief description of the session goals, the current iteration of the template, and summaries of the EU AI Act, NIST AI RMF, and ISO 42001 (Appendix A). We asked participants to familiarize themselves with these materials and optionally provide preliminary feedback by reviewing and annotating the template. This ensured participants understood the key regulatory frameworks, enabling more informed discussions and efficient use of time to revise the template.
@@
Analysis. After each co-design session, two authors conducted an inductive thematic analysis (bottom-up) of the session’s transcripts, following established coding methodologies (Saldaña 2015; Miles and Huberman 1994; McDonald, Schoenebeck, and Forte 2019). The authors used the Figma [^1] https://www.figma.com/ platform to capture participants’ feedback in sticky notes, and collaboratively created themes based on these notes. They discussed and resolved any disagreements that arose during the analysis process to derive a list of usability issues to be addressed in the next iteration of the template.
This co-design sessions with AI practitioners and compliance experts surfaced two additional design requirements, which led to the design of a template in which no further refinements were deemed necessary (Figure 2):
@@
V5 P12 Male 43 PhD machine learning, computer vision 7 researcher
@@
To populate the template, we employed a three-step semi-automatic method, which included soliciting AI practitioners responses and reviewing these responses with AI compliance experts.
In the first step, we compiled a bank of statements to gather responses from AI practitioners. To do so, we sourced statements from responsible AI guidelines (The High-Level Expert Group on Artificial Intelligence 2020; Constantinides et al. 2024a), documentation standards (Selbst 2021; Gebru et al. 2021; Holland et al. 2020; Bender and Friedman 2018; Mitchell et al. 2019; Sokol and Flach 2020; Raji et al. 2020), checklists and impact assessment questionnaires (Madaio et al. 2020; Golpayegani, Pandit, and Lewis 2023; Skoric 2023; National Institute of Standards and Technology 2023b). Next, we reviewed these statements, linking them to the relevant excerpts from the EU AI Act, the NIST AI RMF, and the ISO 42001, and grouping similar ones together. This process resulted in a list of 32 statements grounded in regulations and best responsible AI practices, designed to systematically gather information about the system’s use, components, and data, team involvement, and the associated risks, mitigations, and benefits (Table 3, Appendix C). They also directly map to the sections of our template.
@@
In addition to the report, we included a baseline condition to compare against our final template. For the baseline, we derived a generic impact assessment template for AI system use, based on the typical structure of existing impact assessment reports identified by Stahl et al. (2023). This generic template consisted of three sections: the first, “Intended Use”, included a description of the use (covering the technology and application area) and stakeholders affected by the use; the second,“Risks”, covered human rights risks, ethical risks, data protection and privacy risks, safety risks, and security risks; and the third, “Mitigation measures”, detailed technical and organizational measures for mitigating risks. The baseline content was the same as the final report except for describing intended use through full sentences, reorganizing risks and mitigations under different categories, and lacking the enumeration of benefits. Figures 6 and Figure 7 in Appendix D present the populated final impact assessment report and the baseline used in the user study.
@@
We conducted a user study in the form of semi-structured 30-minute interviews with 8 AI practitioners from the same large tech company and 5 AI compliance experts from both industry and academia. Specifically, our evaluation ought to answer four questions:
@@
Participants. We recruited 8 AI practitioners working in the tech industry, with various roles and expertise, including engineers specializing in computer vision, and researchers in deep-learning. Additionally, 5 AI compliance experts took part in the study, who work in industry (3) and academia (2). Table 2 summarizes participants’ demographics.
@@
Analysis. First, we computed the average Likert scale score for each statement (Figure 4) for both the final report and the baseline. Second, two authors conducted an inductive thematic analysis (bottom-up) of the interview transcripts, following established coding methodologies (Saldaña 2015; Miles and Huberman 1994; McDonald, Schoenebeck, and Forte 2019). The transcripts included how the report’s content provides all information needed for conducting impact assessment, how usable is each template, and any other preferences or dislikes. The authors used the Figma platform to capture participants’ answers in sticky notes, and collaboratively created themes based on these notes. They discussed and resolved any disagreements that arose during the analysis process. Each theme included quotes from at least two participants, signifying that data saturation was reached (Guest, Bunce, and Johnson 2006).
@@
Section 1: Information on the System’s Use and Teams. The section begins with a clear description of the system’s intended use (Figure 3A), including its intended purpose, the technical capabilities that enable this use, the area or sector chosen for its application (including its geographic and temporal extent), any natural persons and groups likely to be affected by this use, and any natural or legal persons who have authority over it. This description mirrors the five risk assessment components identified by Golpayegani (Goodell et al. 2021): purpose, capability, domain, user, and subject that provide information necessary for conducting the risk assessment based on the EU AI Act (addressing R1). Users could specify these components using external dictionaries, such as the Vocabulary of AI Risks (Golpaygani, Pandit, and Lewis 2023), which contains pre-defined descriptions for each of these concepts sourced directly from the Act.
@@
Section 2: Risks. This section lists potential risks (Figure 3B) associated with putting the system into use (addressing R2). To begin with, it presents a brief summary statement outlining the main risks that may undermine safety, rule of law, fundamental rights, health, environment, and democracy. It then lists all the specific risks of system’s use grouped into risks stemming from the system’s capability (i.e., the technical components of the AI system), human interaction (i.e., experiences of people interacting with the AI system), and systemic impact (i.e., societal, economic, and environmental impacts of the AI system’s use). This grouping aligns with a three-layered framework for evaluating sociotechnical harms by Google DeepMind (Weidinger et al. 2023). It was selected because AI practitioners found it more approachable (addressing R4), based on feedback indicating difficulties in envisioning realistic risks across stakeholders (i.e., organizations, individuals, groups of individuals, and societies) listed in the EU AI Act, AI RMF, and ISO 42001. For each risk listed in the section, we have added to the template two empty placeholders specifically for ‘Risk likelihood’ and ‘Risk level’. This encourages users of the template to categorize each risk’s impact by estimating its probability and potential impact as high, medium, low, respectively, and record these assessments in the provided placeholders. This facilitates impact assessment per ISO 42001 (addressing R1) and allows estimation of the overall risk level score of the use as a combination of the risk likelihoods and risk levels.
@@
P2 Male 30 PhD machine learning 5 researcher
@@
P4 Male 32 PhD machine learning, NLP 7 researcher
P5 Male 35 PhD embedded machine learning 5 manager
P6 Male 28 PhD mobile sensing 6 researcher

### v2

@@ Abstract
footnotetext: Project website: https://social-dynamics.net/impact-assessment
@@ Introduction
As it has become evident that AI system impacts involve not only the underlying algorithms (Watkins et al. 2021; Barnett and Diakopoulos 2022) but also their systemic impacts (Ehsan et al. 2022; Shelby et al. 2023), there has been a growing demand for comprehensive AI impact assessments (AIIAs) (Selbst 2021). In the absence of regulatory requirements and standardized frameworks, initially, AIIAs have been regarded as an extension of existing AI governance processes, such as assessments of privacy (Wright and Wadhwa 2012), data protection (Janssen 2020), and social consequences of AI systems, models and services (Selbst 2021; Raji et al. 2020). However, over time, AIIAs developed into an independent component of AI governance (Skoric 2023). By 2021, a total of 38 distinct impact assessment methods have been introduced (Stahl et al. 2023) serving various goals such as prompting companies to proactively address social consequences (Sherman and Eisenberg 2024), aligning system behavior with organization’s responsible AI principles (Raji et al. 2020; Fujitsu Research Center for AI Ethics 2022; Microsoft 2022b), increasing awareness of potential harms among development teams (Johnson and Heidari 2023), and documenting decisions to facilitate learning and future governance development (Selbst 2021).
However, current AIIAs reports are marked by a lack of comprehensiveness and insufficient grounding in established frameworks such as the EU AI Act (European Comission 2024), the NIST’s AI Risk Management Framework (AI RMF) (National Institute of Standards and Technology 2023a), and international standards such as the ISO 42001 AI Management System (ISO/IEC 2023), that provide complementary angles for impact assessment: regulatory compliance (EU AI Act), risk management (AI RMF), and organizational best practices (ISO 42001). The lack of alignment with regulations makes it harder for stakeholders to understand the broad societal, environmental, and economic impacts of AI uses, which are essential for risk assessments (European Comission 2024; Hupont et al. 2024; Stahl et al. 2023). Instead, they must repeatedly gather information across disparate assessments of various AI components (e.g., models, systems, and algorithms) rather than focusing on specific uses. Developer teams, in particular, often encounter difficulties in initiating AI impact assessments (Buçinca et al. 2023) and require additional guidance throughout this process (Wang et al. 2023).
In response to these issues, our aim is to examine recent regulatory frameworks and existing AIIAs to devise together with AI practitioners and AI compliance experts a reporting template focusing on the intended use of the system. This template aims to provide the necessary information for conducting impact assessments based on regulations across various AI system uses and to be adaptable to different roles. In so doing, we made two contributions:
@@
    We designed a comprehensive template for an impact assessment report that is grounded in regulations, and did so by conducting two studies. First, we elicited an initial set of design requirements through literature review and semi-structured interviews with 2 AI compliance experts. Second, we co-designed the impact assessment report template with 14 AI practitioners and 6 AI compliance experts, grounded it in regulatory requirements of the EU AI Act, AI RMF, and ISO 42001.
@@
    We populated the template with a real-world use case of an AI system (a companion app designed to improve the meeting experience) and evaluated it with 8 AI practitioners and 5 AI compliance experts from academia and industry. We compared the final report to a baseline report derived from the typical structure of existing impact assessment reports. Participants found that the final report provided more complete information for impact assessments and addressed all AI system components and impacts more broadly than the baseline. Both the final report and the baseline were rated as highly adaptable to different AI uses and adaptable to different roles.
@@
Figure 1: Overview of our four-step method for designing a comprehensive template for an impact assessment report grounded in the EU AI Act, NIST’s AI RMF, and ISO 42001. In the first step, we interviewed AI compliance experts to elicit two design requirements and design the initial prototype of the template (V1). In the second step, we engaged both AI practitioners and AI experts to iteratively elicit two additional design requirements and co-design four iterations of the template (V2-V5). In the third step, we populated the final version of the template with AI practitioners’ responses. In the fourth step, we evaluated it against our four design requirements and the baseline template identified in the literature review of AIIAs (Stahl et al. 2023).
@@
AI impact assessments are defined as structured processes for understanding the implications of proposed AI systems (Stahl et al. 2023). They have been proposed as both law-agnostic self-regulating processes (National Institute of Standards and Technology 2023a, b) and official processes in regulations (European Comission 2024; The Danish Institute for Human Rights 2023; The Government of Canada 2023) and organizational standards (ISO/IEC 2023). For example, the NIST AI RMF, a voluntary guidance framework for organizations that design, develop, deploy, or use AI systems, suggests assessing the beneficial and harmful impacts of AI on individuals, groups, organizations, and society. The EU AI Act (European Comission 2024), the first comprehensive AI regulation in the European Union, requires fundamental rights impact assessment for high-risk AI uses deployed by bodies governed by public law, private operators providing public services, and operators assessing creditworthiness and conducting risk assessments for life and health insurance. This impact assessment, conducted before deployment and updated as needed, should cover affected groups, risks of harm, human oversight, and risk management strategies. Similarly, the world’s first ISO 42001 standard for AI management system (ISO/IEC 2023) requires organizations to assess the potential consequences of AI systems on individuals and societies, including human rights, legal positions, and life opportunities.
Despite ongoing efforts to standardize AIIAs (ISO/IEC 2025), a consensus on their content and reporting methods for stakeholders and the public is still lacking (Watkins et al. 2021; Sherman and Eisenberg 2024). Stahl et al. (2023) identified and studied 38 AIIAs, showing that they vary widely in topics, focus, and formats, primarily covering human rights (The Council of Europe 2018; Mantelero 2022) and ethics (Gebru et al. 2021). The focus of AIIAs varies, assessing the impacts of AI models, systems, or services, typically mentioning use cases only briefly in final reports. However, there is growing agreement that impact assessments should be conducted on specific uses (European Comission 2024; Hupont et al. 2024). In the EU AI Act (European Comission 2024), the specific use determines one of four risk categories (unacceptable, high, limited, minimal), each with its own legal requirements. Similarly, in the NIST framework (National Institute of Standards and Technology 2023a), use-case profiles are required to describe the current state and the desired state of the system, allowing for risk management at different stages of the AI lifecycle.
@@
Researchers highlight the need for more inclusive, value-sensitive co-design processes (Johnson and Heidari 2023; Sadek et al. 2024) that explicitly incorporate compliance experts (Hupont et al. 2024; Constantinides et al. 2024a). For example, OpenLoop’s policy prototyping experiment involved developers, researchers, policymakers, and regulators collaborating to create a new policy and documentation of potential harms and mitigations for automated decision-making systems (Gomes de Andrade and Kontschieder 2021). Similarly, the creators of use case cards detailing AI systems’ intended uses engaged in co-design workshops with Unified Modeling Language (UML) experts, and EU policy experts (Hupont et al. 2024).
@@ Methods
To design a comprehensive template for an impact assessment report that is grounded in regulations, we conducted a series of four studies (Figure 1). These studies included literature review and interviews with 2 AI compliance experts, an iterative co-design process with 14 AI practitioners and 6 AI compliance experts, review of the populated template, and a user study to evaluate the populated template with 8 AI practitioners working in industry research and 5 AI compliance experts from industry and academia.
@@
To elicit requirements for designing a comprehensive template for an impact assessment report that is grounded in regulations, we resorted to previous literature and did so in two steps. First, we reviewed prior academic literature investigating the use of risk assessments, impact reports, and AI documentation for compliance (Skoric 2023; Selbst 2021; Gaumond and Régis 2023; National Institute of Standards and Technology 2023a). We also drew upon five existing AI impact assessment templates (Microsoft 2022b; Fujitsu Research Center for AI Ethics 2022; National Institute of Standards and Technology 2023b; Ada Lovelace Institute 2022; The Government of Canada 2023) and value-sensitive games and cards (Ballard, Chappell, and Kennedy 2019; Vakkuri et al. 2021; Sadek et al. 2024). Second, we identified the necessary minimum information to conduct impact assessment in line with three frameworks: the EU AI Act (European Comission 2024), the NIST’s AI RMF (National Institute of Standards and Technology 2023a), and the ISO 42001 (ISO/IEC 2023). We selected these frameworks because they have undergone a consensus-driven, transparent process with wide consultations involving various roles such as developers, researchers, managers, and compliance experts. Together, they provide three complementary angles for impact assessment: the EU AI Act offers regulatory compliance, setting a precedent for other AI regulations; the AI RMF provides practical insights for risk management in corporate environments; and ISO 42001 emphasizes organizational best practices. In this step, we started from reviewing the EU AI Act (European Comission 2024) and identifying relevant excerpts that pertain to risk management (Article 9), data governance (Article 10), system monitoring (Articles 14, 15, 72), technical documentation (Articles 11, 12, 18, Annex IV), system transparency (Articles 13, 50), impact assessment (Article 27), and provider obligations (Articles 16, 17, 53, 55). We then reviewed the NIST’s AI RMF (National Institute of Standards and Technology 2023a) and its Algorithmic Impact Assessment template (National Institute of Standards and Technology 2023b). From these two sources, we identified relevant excerpts related to mapping the risks of a system’s use, minimizing its negative impacts while maximizing its benefits, and communicating these aspects to various stakeholders. Finally, from the ISO 42001, we identified excerpts on documenting resources for AI systems (data, tools, systems, computing, and human resources), assessing impacts on individuals, groups of individuals, and society, and reporting AI system details to stakeholders.
@@
Having these excerpts from previous literature at hand, we conducted 30-minute semi-structured interviews with 2 AI compliance experts to jointly annotate the collected excerpts (i.e., the EU AI Act, the NIST’s AI RMF, and the ISO 42001) with high-level topics. We then jointly organized the topics in a thematic and hierarchical manner, resulting in the creation of the first version of the template (Figure 1, Step 1, V1). Through this joint exercise, we surfaced two main design requirements for an impact assessment report template:
@@
After designing the first version of the template, we conducted a series of individual, 30-minute co-design sessions with 14 AI practitioners and 6 AI compliance experts (Figure 1, Step 2). These sessions aimed to ensure the template provided sufficient information for impact assessment per the EU AI Act, NIST’s AI RMF, and ISO 42001, and was usable by various roles in the assessment process.
Participants. We aimed to achieve a diverse participant sample using snowball sampling, a method where existing study participants recruit future participants from among their acquaintances. We began by identifying 6 initial participants through an internal mailing list at a large tech company, seeking individuals familiar with the EU AI Act and actively involved in developing or evaluating at least one ongoing AI project during the time of the planned interviews. These participants were then asked to refer additional participants from their networks, thus expanding the sample size through successive referrals. We recruited a total of 20 participants (13 male, 7 female, with a median age of 34 years old) representing a variety of AI practitioners such as researchers (9), engineers (2), managers (3), designers (1), and AI compliance experts (6), and potential end users of the template. Their expertise span across various areas including generative AI, deep learning, AI standardization, and human rights. AI practitioners offered practical examples on navigating regulatory challenges in their projects while maintaining innovation. Compliance experts contributed their knowledge of existing regulations, insights from ongoing regulatory works, and lessons learned from reviewing AI system uses. Table 1 provides an overview of the participants’ demographics and relevant experience.
Procedure. Before the session, we emailed participants with a demographic survey, a brief description of the session goals, the current iteration of the template, and summaries of the EU AI Act, NIST AI RMF, and ISO 42001 (Appendix A). We asked participants to familiarize themselves with these materials and optionally provide preliminary feedback by reviewing and annotating the template. This ensured participants understood the key regulatory frameworks, leading to more informed discussions and template revisions.
@@
Analysis. After each co-design session, two authors conducted an inductive thematic analysis (bottom-up) of the session’s transcripts, following established coding methodologies (Saldaña 2015; Miles and Huberman 1994; McDonald, Schoenebeck, and Forte 2019). The authors used the Figma platform (Figma 2024) to capture participants’ feedback in sticky notes, and collaboratively created themes based on these notes. They discussed and resolved any disagreements that arose during the analysis to derive a list of usability issues to be addressed in the next iteration of the template.
These co-design sessions with AI practitioners and compliance experts surfaced two additional design requirements, which led to the design of a template in which no further refinements were deemed necessary (Figure 1, Step 2, V5):
@@
V5 P12 Male 43 PhD machine learning, computer vision 5 researcher
@@
To populate the template, we employed a three-step semi-automatic method, which included soliciting AI practitioners responses and reviewing these responses with AI compliance experts (Figure 1, Step 3).
In the first step, we compiled a list of statements to gather responses from AI practitioners. To do so, we sourced statements from responsible AI guidelines (The High-Level Expert Group on Artificial Intelligence 2020; Constantinides et al. 2024a), documentation standards (Selbst 2021; Gebru et al. 2021; Holland et al. 2020; Bender and Friedman 2018; Mitchell et al. 2019; Sokol and Flach 2020; Raji et al. 2020), checklists and impact assessment questionnaires (Madaio et al. 2020; Golpayegani, Pandit, and Lewis 2023; Skoric 2023; National Institute of Standards and Technology 2023b). Next, we reviewed these statements, linking them to the relevant excerpts from the EU AI Act, the NIST AI RMF, and the ISO 42001, and grouping similar ones together. This process resulted in a list of 32 statements grounded in regulations and best responsible AI practices, designed to systematically gather information about the system’s use, components, and data, team involvement, and the associated risks, mitigations, and benefits (Appendix C). They also directly map to the sections of our template.
@@
In addition to the report, we included a baseline condition to compare against our final template. For the baseline, we derived a generic impact assessment template for AI system use, based on the typical structure of existing impact assessment reports identified by Stahl et al. (2023). This generic template consisted of three sections: the first, “Intended Use”, included a description of the use (covering the technology and application area) and stakeholders affected by the use; the second,“Risks”, covered human rights risks, ethical risks, data protection and privacy risks, safety risks, security risks, and environmental risks; and the third, “Mitigation measures”, detailed technical and organizational measures for mitigating risks.
The baseline content was the same as the final report except for describing intended use through full sentences, reorganizing risks and mitigations under different categories, and lacking the enumeration of benefits. Figures 5 and Figure 6 in Appendix D present the populated final impact assessment report and the baseline used in the user study.
@@
We conducted a user study in the form of semi-structured 30-minute interviews with 8 AI practitioners from the same large tech company and 5 AI compliance experts from both industry and academia (Figure 1, Step 4). Specifically, our evaluation ought to answer four questions:
@@
Participants. We recruited 8 new AI practitioners working in the tech industry, with various roles and expertise, including computer vision engineers and researchers in deep learning, different from those who participated in the co-design sessions. Additionally, 5 new AI compliance experts took part in the study, who work in industry (3) and academia (2). Table 2 summarizes participants’ demographics.
@@
Analysis. First, we computed the average Likert scale score for each statement (Figure 3) for both the final report and the baseline. Second, two authors conducted an inductive thematic analysis (bottom-up) of the interview transcripts, following established coding methodologies (Saldaña 2015; Miles and Huberman 1994; McDonald, Schoenebeck, and Forte 2019). The transcripts covered how the report’s content supports impact assessment, how usable each template is, and any other preferences or dislikes.
The authors used the Figma platform (Figma 2024) to capture participants’ answers in sticky notes, and collaboratively created themes based on these notes. They discussed and resolved any disagreements that arose during the analysis process. Each theme included quotes from at least two participants, signifying that data saturation was reached (Guest, Bunce, and Johnson 2006).
@@
Section 1: Information on the System’s Use and Teams. The section begins with a clear description of the system’s intended use (Figure 2A), including its intended purpose, the technical capabilities that enable this use, the area or sector chosen for its application (including its geographic and temporal extent), any natural persons and groups likely to be affected by this use, and any natural or legal persons who have authority over it. This description mirrors the five risk assessment components identified by Golpayegani, Pandit, and Lewis (2023): purpose, capability, domain, user, and subject that provide information necessary for conducting the risk assessment based on the EU AI Act (addressing R1). Users could specify these components using external dictionaries, such as the Vocabulary of AI Risks (Golpaygani, Pandit, and Lewis 2023), which contains pre-defined descriptions of these concepts directly sourced from the Act.
@@
Section 2: Risks. This section lists potential risks (Figure 2B) associated with putting the system into use (addressing R2). To begin with, it presents a brief summary statement outlining the main risks that may undermine safety, rule of law, fundamental rights, health, environment, and democracy. It then lists all the specific risks of system’s use grouped into risks stemming from the system’s capability (i.e., the technical components of the AI system), human interaction (i.e., experiences of people interacting with the AI system), and systemic impact (i.e., societal, economic, and environmental impacts of the AI system’s use). This grouping aligns with a three-layered framework for evaluating sociotechnical harms by Google DeepMind (Weidinger et al. 2023). It was selected because AI practitioners found it more approachable (addressing R4), based on feedback indicating difficulties in envisioning realistic risks across stakeholders (i.e., organizations, individuals, groups of individuals, and societies) listed in the EU AI Act, AI RMF, and ISO 42001. For each risk listed in the section, we have added to the template two placeholders specifically for ‘Risk likelihood’ and ‘Risk level’. This encourages users of the template to categorize each risk’s impact by estimating its probability and potential impact as high, medium, low, respectively, and record these assessments in the provided placeholders. This facilitates impact assessment per ISO 42001 (addressing R1) and allows estimation of the overall risk level score of the use as a combination of the risk likelihoods and risk levels.
@@
P2 Male 31 PhD machine learning 5 researcher
@@
P4 Male 32 PhD machine learning 7 researcher
P5 Male 36 PhD embedded machine learning 5 manager
P6 Male 27 PhD mobile sensing 6 researcher